OpenStudy (anonymous):

The measure of the supplement of an angle exceeds three times the measure of the complement of the angle by 12. Find the measure of half of the supplement -please show steps- The measure of the supplement of an angle exceeds three times the measure of the complement of the angle by 12. Find the measure of half of the supplement -please show steps- @Mathematics

OpenStudy (anonymous):

180-x=3(90-x)+12 solve for x the supplement is 180-x

OpenStudy (anonymous):

i got that part. but then x=-102 and if you plug that in, is 282 which is more then 180, which doesn't make sense

OpenStudy (anonymous):

180-x=270-3x+12 2x=102 x=51??

OpenStudy (anonymous):

o haha oops

OpenStudy (anonymous):

wait so then half the supplement would be 51/2?

OpenStudy (anonymous):

the supplement is 180-51=129 and half of it is 64.5

OpenStudy (anonymous):

um x=25.5 not 51

OpenStudy (anonymous):

how if 2x=102 then x is 51 but x is not the supplement ,the supplement is 180-x

OpenStudy (anonymous):

it's -4x=102

OpenStudy (mertsj):

x is the angle. 180-x is the supplement. 90 - x is the complement. Three times the complement is 3(90-x). The problem says the supplement is bigger than 3 times the complement. So if you want an equation, you must either add to the little one or subtract from the big one. Since the supplement is 12 bigger, let's subtract 12 from it: supplement less 12 is equal to 3 times the complement (180 -x ) -12 = 3(90-x) Now solve that: 180-x-12 = 270 - 3x 168 - x = 270 - 3x Add 3x 168 + 2x = 270 Subtract 168 2x = 102 Divide by 2 x = 52 x is the angle. the supplement is 180-x or 180-52 or 128 So half the supplement is 1/2 of 128 or 64

OpenStudy (anonymous):

why is -12 on the left side? you didn't need to subtract it. You were supposed combine like terms on the right side making it 282

OpenStudy (mertsj):

Because the supplement is bigger. If i want them to be equal, I can subtract from the supplement or add to 3 times the complement. I chose to subtract from the supplement.

OpenStudy (mertsj):

There are no like terms on the right to combine.

OpenStudy (anonymous):

yes there are. 270 and 12 are like terms

OpenStudy (mertsj):

If you want to add to the little one which is 3 times the complement, then the equation would be: 180-x = 3(90-x) +12 180 - x = 270 -3x +12 180 - x = 282 - 3x Add x to both sides: 180 = 282 -2x Subtract 282 from both sides: -102 = -2x Divide by -2: 52 = x So the angle is 52 degrees. The supplement is 180 - 52 or 128 degrees One half the supplement is 1/2 of 128 or 64 degrees.

OpenStudy (mertsj):

Did you even look at the first solution I gave you? There were NO like terms on the right.

OpenStudy (anonymous):

-3-1 is -4 not -2

OpenStudy (mertsj):

It says to ADD x to both sides -3x + 1x = -2x

OpenStudy (anonymous):

ok i did 180-x=3(90-x)+12 180-x=270+3x+12 180-x=282+3x -180 -180 -x=102+3x -3x -3x -4x=102 /-4 /-4 x=-25.5

OpenStudy (mertsj):

3(90-x) = 270 - 3x NOT 270 +3x

OpenStudy (anonymous):

oooooooooooooooooo. omg that was so stupid of me!!!

OpenStudy (mertsj):

Wake up and die right

OpenStudy (anonymous):

die right? what is that?

OpenStudy (mertsj):

Just an old saying. But what I don't understand is why you did not look at either one of the two correct solutions I gave you. You would have found that error.

OpenStudy (anonymous):

i did look i guess i just didn't notice

OpenStudy (anonymous):

wait. ok so then you get 102/2= 51

OpenStudy (anonymous):

not 52

OpenStudy (mertsj):

Yep. So the angle is 51 and the supplement is 129 and half the supplement is 64.5

OpenStudy (anonymous):

ooooo. THANKS SO MUCH!!!

OpenStudy (mertsj):

So the angle is 51, the supplement is 129 and the complement is 39. Three times the complement is 3(39) or 117. add twelve to it and you get the supplement which is 129.